# ContractSkill: Repairable Contract-Based Skills for Multimodal Web Agents

### 20 Mar 2026 v1

@@ Abstract.
Despite rapid progress in multimodal GUI agents, reusable skill acquisition remains difficult because on-demand generated skills often leave action semantics, state assumptions, and success criteria implicit. This makes them brittle to execution errors, hard to verify, and difficult to repair. We present ContractSkill, a framework that converts a draft skill into a contracted executable artifact with explicit preconditions, step specifications, postconditions, recovery rules, and termination checks. This representation enables deterministic verification, step-level fault localization, and minimal patch-based repair, turning skill refinement into localized editing rather than full regeneration. Experiments on VisualWebArena and MiniWoB with GLM-4.6V and Qwen3.5-Plus show that ContractSkill improves self-generated skills from 9.4% and 10.9% to 28.1% and 37.5% on VisualWebArena, and from 66.5% and 60.5% to 77.5% and 81.0% on MiniWoB. Repaired artifacts also transfer across models, improving the target model’s self-generated-skill baseline by up to 47.8 points and 12.8 points on the two benchmarks, respectively. These results suggest that agent skills are better treated as explicit procedural artifacts that can be verified, repaired, and shared across models.
multimodal web agents, skill, deterministic verifier, program repair, cross-model transfer
$c$s: Computing methodologies Intelligent agentsccs: Human-centered computing Graphical user interfacesccs: Software and its engineering Automatic programming
@@ 1. Introduction
Multimodal web agents now operate in increasingly realistic browser environments, where a model must jointly interpret screenshots, structured page information, and natural-language goals while choosing actions across long interaction horizons. VisualWebArena has made this setting concrete and measurable for visually grounded web interaction, and it highlights both the promise and the brittleness of current systems (Koh et al., 2024). The problem is inherently multimedia-centric because success depends on coordinating visual grounding, interface semantics, and sequential decision making in a shared action loop.
To address long-horizon execution, many recent agents rely on self-generated procedural guidance. In mobile and web settings, systems often produce a textual skill, manual, or reflection trace before or during action execution (Zhang et al., 2025; Wang et al., 2024b; Chen et al., 2024). These artifacts help, but they are still awkward knowledge objects. They are usually generated on demand, tailored to a single run, and expressed as loosely structured text. As a result, they behave less like durable skills and more like one-off procedural drafts.
Recent benchmark evidence reinforces this concern. SkillsBench reports that self-generated skills provide negligible or negative benefit on average across diverse tasks, which suggests that current models do not yet reliably author high-quality procedural knowledge for later reuse (Li et al., 2026).
This creates a practical failure mode. Self-generated skills usually do not specify what must hold before a step, what concrete page evidence should hold after the step, or what recovery policy should be applied when the page state deviates from expectation. Once execution fails, the system often falls back to rewriting the entire skill description or regenerating a new trajectory. This strategy may repair the immediate run, but it does not give the agent a clean, externally maintainable knowledge object.
Existing methods only partially solve this problem. Reflection-based or self-repair methods primarily update the current episode. They revise action choices, reasoning traces, or transient plans based on observed feedback (Shinn et al., 2023; Wang et al., 2024b). Skill-bank style work, in contrast, focuses on larger repositories of reusable knowledge, often in open-ended environments (Wang et al., 2024a; Chen et al., 2025). Our target is different from both. We focus on a single explicit skill artifact for multimodal web interaction and ask how to make it executable, verifiable, locally repairable, and consumable by another model.
We therefore propose ContractSkill, a contract-based framework for repairing skill artifacts in multimodal web agents. Starting from a model-generated draft skill, ContractSkill compiles the draft into an executable artifact with preconditions, step objects, postconditions, recovery policies, and termination checks. A deterministic verifier then evaluates each step against page-state evidence and returns a structured error code together with the failing step index. Instead of rewriting the whole skill, ContractSkill applies one or a few minimal patch operators that change only the relevant selector, precondition, postcondition, recovery rule, or action argument. The repaired artifact can be reused by the source model and, importantly, consumed by a target model that did not generate the skill in the first place.
@@
  • We introduce a contract-based representation that compiles a self-generated textual skill into a repairable skill artifact with explicit execution semantics.
  • We design a verifier-guided repair framework that combines deterministic page-state checking, step-level fault localization, and minimal patch operators.
  • We formulate a cross-model skill transfer setting in which a repaired artifact produced by one model can be executed by another, showing that repaired AI-generated skills can become portable procedural assets rather than model-specific prompts.
@@ 2. Related Work
2.1. Multimodal Web and GUI Agents
Web-agent benchmarks span a spectrum from controlled browser tasks to realistic multimodal web workflows. Early environments such as World of Bits and MiniWoB established reproducible settings for studying fundamental web interaction skills, while later platforms such as WebArena moved toward longer-horizon and more realistic web execution (Shi et al., 2017; Liu et al., 2018; Zhou et al., 2024). VisualWebArena further extends this line to visually grounded web tasks that require agents to interpret image-text inputs and act on realistic websites (Koh et al., 2024). Together, MiniWoB and VisualWebArena provide complementary evaluation settings, with MiniWoB emphasizing controlled and repeatable interaction patterns and VisualWebArena emphasizing realistic multimodal web interaction.
@@ 2.2. Skill Artifacts and Procedural Knowledge
Recent work increasingly treats agent skills as explicit and portable procedural knowledge rather than transient prompting context. Agent Skills for Large Language Models surveys this abstraction layer, SkillsBench argues that skill quality and transferability should be evaluated directly, and analysis of single-agent skill libraries studies when external skills reduce coordination cost and when they fail (Xu and Yan, 2026; Li et al., 2026; Li, 2026).
Closer to our setting, Agent Workflow Memory stores reusable textual workflows, ASI induces and verifies executable programmatic skills online, and PolySkill separates abstract goals from concrete implementations for cross-site reuse (Wang et al., 2024c, 2025; Yu et al., 2025). SkillWeaver discovers website-specific APIs through iterative self-improvement, Hierarchical Memory Tree organizes web experience with intent-, stage-, and action-level memory plus explicit pre/post conditions, and RAG-GUI retrieves external GUI guidelines at inference time (Zheng et al., 2025; Tan et al., 2026; Xu et al., 2025). These works reinforce the value of structured procedural knowledge, but our focus is narrower. We study a single externalized multimodal web skill artifact with deterministic step-level diagnosis, minimal post-hoc repair, and cross-model reuse.
2.3. Self-Repair and Reflection in Agents
Recent work on self-improvement and reflection studies how agents iteratively refine behavior through memory, reflection, and evolution. ParamMem augments language agents with parametric reflective memory to increase the diversity and usefulness of self-reflection signals across tasks (Yao et al., 2026). Group-Evolving Agents explores open-ended self-improvement through experience sharing and repeated evolution at the group level (Weng et al., 2026). Despite these advances, the target of refinement in this line of work remains primarily the agent’s current behavior, trajectory, or internal update process. By contrast, our work repairs the externalized skill artifact itself. This difference matters because the repaired artifact can persist as a reusable object across episodes and can also be consumed by another model.
@@ 2.4. Program Repair and Verifier-Guided Optimization
Program repair offers a useful methodological lens for our setting because it emphasizes three principles that are also central here, namely localizing failures, constraining the edit space, and using a verifier as the acceptance oracle. Classic automatic program repair spans generate-and-validate, template-based, semantics-based, and learning-guided methods, as exemplified by GenProg, PAR, SemFix, DirectFix, Angelix, and Prophet (Goues et al., 2012; Kim et al., 2013; Nguyen et al., 2013; Mechtaev et al., 2015, 2016; Long and Rinard, 2016). More broadly, verifier-guided refinement frameworks such as counterexample-guided abstraction refinement show how candidate solutions can be iteratively falsified and improved through structured feedback (Clarke et al., 2000). Although our setting does not repair source code directly, ContractSkill follows the same discipline. It treats a skill as a program-like artifact with contracts, uses deterministic checks to identify the failing component, and applies minimal local edits rather than unconstrained rewrites. This perspective brings verifier-guided repair principles into multimodal web agent skills.
@@ 3.1. Environment
We consider a multimodal web environment, following recent execution-based agent benchmarks such as VisualWebArena, WorkArena, and OSWorld (Koh et al., 2024; Drouin et al., 2024; Xie et al., 2024), in which an agent receives a natural-language task instruction $x$. At step $t$, the environment provides an observation $o_{t}=(I_{t},D_{t})$, where $I_{t}$ is the current page screenshot and $D_{t}$ is a structured page summary, such as a DOM abstraction, accessibility summary, or textual page state description. The agent outputs an action $a_{t}\in\mathcal{A}$, where $\mathcal{A}$ includes operations such as click, type, scroll, and select. The environment then transitions to a new observation $o_{t+1}$.
@@ 4.1. Skill Artifact Schema
ContractSkill represents each skill as an explicit artifact with the following fields, namely skill_name, goal, preconditions, steps, postconditions, recovery, and terminate. Each step has four mandatory elements, namely selector, action, args, and post_assertion. Compared with free-form text, this schema exposes the exact operational assumptions behind the skill. The schema is designed to externalize procedures as reusable objects, which is also the motivating direction behind manual construction and page-level knowledge extraction in recent agent systems.
@@ 4.2. Deterministic Verifier
where $i$ is the failed step index, $e$ is a structured error code, and $\tau_{i}$ is the local trace evidence used for diagnosis. Verification uses deterministic checks such as URL matching, DOM retrieval, textual assertions, form-value checks, and state-change predicates. In practice, selectors are resolved on the current accessibility or DOM snapshot using role, visible text, and stable attributes rather than persistent node ids. Postconditions are instantiated from action-specific templates, such as value matching for type, state or navigation change for click, and option validation for select. This design keeps the repair signal grounded in observable page state rather than in model self-evaluation, matching the execution-oriented evaluation logic used in WebArena, VisualWebArena, and WorkArena (Zhou et al., 2024; Koh et al., 2024; Drouin et al., 2024).
@@ 4.4. Minimal Patch Operators
Each operator changes only the artifact fragment tied to the diagnosed failure. This design preserves the rest of the artifact and avoids the instability that often accompanies full-text rewrites. The operator set intentionally favors small, interpretable edits, echoing the minimality bias emphasized in DirectFix and learning-guided repair systems such as Prophet (Mechtaev et al., 2015; Long and Rinard, 2016).
@@ 4.5. Verifier-Guided Repair Loop
(9) $s^{\prime}\text{ is accepted if }J(s^{\prime})>J(s).$
@@ 5. Experimental Setup
Our experiments are organized around three questions. We ask (1) whether ContractSkill improves end-to-end web execution over direct acting and naive skill use, (2) whether any improvement comes specifically from verifier-guided local repair rather than rewriting alone, and (3) whether repaired skill artifacts can transfer across models. We structure the evaluation accordingly.
@@ 5.3. Baselines
We compare against two core baselines in all main-result tables. No-Skill executes the task directly without any external procedural guidance, representing the standard direct-acting setting. Self-Generated Skill lets the acting model produce a skill and immediately use it, but does not perform verifier-guided repair after failure. This comparison separates the value of skill writing itself from the value of repair. On MiniWoB, we additionally include more controlled comparisons. Text-Only Rewrite rewrites the skill after failure in free-form text, without contracts, structured error codes, or localized edits, and therefore isolates the effect of rewriting alone. We also report two ablations of ContractSkill. No Failure Localization removes targeted identification of the failing skill component, and Unconstrained Repair permits broad revisions rather than restricted local fixes. These controlled variants help determine whether the gains of ContractSkill arise from structured verifier-guided repair.
@@ 5.4. Metrics
Our primary metric is verifier-based success rate. To characterize efficiency, we also report average steps and average total tokens whenever those fields are present in the cleaned summaries. We interpret efficiency as a success-cost trade-off rather than strict dominance over every baseline, because a repaired artifact may spend extra steps or tokens to complete tasks that a brittle self-generated skill fails to finish. Tokens are reported as average total tokens per task and abbreviated in thousands (k). For transfer, we compare repaired-artifact execution against the target model’s own self-generated skill baseline and report both absolute success and the point improvement.
@@ 6. Main Results
We report the main results in the order of argumentative strength. We first test whether ContractSkill works in realistic web interaction, then examine the same claim in a controlled environment where the mechanism is easier to isolate, and finally test whether the repaired artifact transfers across models.
@@ 6.1. VisualWebArena
Table 3 reports the main VWA results. The overall pattern is clear. On realistic web tasks, simply having a skill is not enough, but repairing it into a contracted artifact is consistently beneficial. With GLM-4.6V, ContractSkill reaches 28.1% success, compared with 12.5% for No-Skill and 9.4% for Self-Generated Skill. The efficiency profile also improves. Average steps drop from 12.59 for No-Skill to 3.62, and average token usage decreases from 52.7k to 46.45k. With Qwen3.5-Plus, the packaged final verified aggregation raises ContractSkill to 37.5%, exceeding both No-Skill (28.1%) and Self-Generated Skill (10.9%), while keeping average token usage slightly below No-Skill (40.1k vs. 42.6k). These results matter because VWA is precisely the regime in which brittle skills tend to fail, with long trajectories, changing page state, and multimodal ambiguity. The fact that ContractSkill improves both models suggests that the gain comes from structured repair of the external artifact rather than from model-specific prompting tricks.

### 15 Apr 2026 v3

@@ Abstract.
Self-generated skills for web agents are often unstable and can even hurt performance relative to direct acting. We argue that the key bottleneck is not only skill generation quality, but the fact that web skills remain implicit and therefore cannot be checked or locally repaired. To address this, we present ContractSkill, a framework that converts a draft skill into an executable artifact with explicit procedural structure, enabling deterministic verification, fault localization, and minimal local repair. This turns skill refinement from full rewriting into localized editing of a single skill artifact. Experiments on VisualWebArena show that ContractSkill is effective in realistic web environments, while MiniWoB provides a controlled test of the mechanism behind the gain. Under matched transfer layers, repaired artifacts also remain reusable after removing the source model from the loop, providing evidence of portability within the same benchmark family rather than full-benchmark generalization. These results suggest that the central challenge is not merely generating skills, but making them explicit, executable, and repairable. Code is available at https://github.com/underfitting-lu/contractskill.git.
Skill, Multimodal Web Agents, Contracted Executable Artifacts, Self-Generated, Cross-Model Transfer
$c$s: Computing methodologies Intelligent agentsccs: Human-centered computing Empirical studies in HCIccs: Information systems World Wide Web
@@ 1. Introduction
Long-horizon web tasks seem like exactly the setting where skills should help most. Yet recent evidence already points to a striking mismatch between that intuition and current practice. SkillsBench (Li et al., 2026) reports that curated skills raise average pass rate by 16.2 percentage points, whereas self-generated skills provide no benefit on average . For us, this is the real motivating anomaly. It suggests that the problem is not whether external procedural knowledge can help agents, but whether model-written skills are stable enough to serve as executable objects in the first place.
Our diagnosis is object-centered rather than pipeline-centered. The key bottleneck is not only skill generation quality, but the fact that web skills remain implicit and therefore cannot be checked or locally repaired. Most draft skills are loose textual objects: they omit explicit preconditions, blur step boundaries, leave success evidence underspecified, and provide no principled recovery policy when the page state diverges. Once execution fails, the agent typically rewrites the whole skill or abandons it altogether.
This diagnosis also clarifies the gap with prior work. Reflection and self-repair methods update trajectories, memories, or transient plans in the current episode (Shinn et al., 2023; Wang et al., 2024b). Skill-bank style work maintains larger workflow collections or reusable knowledge repositories in open-ended settings (Wang et al., 2024a; Chen et al., 2025). Our target is different. We update the single skill artifact itself. The question we ask is whether a model-written web skill can be turned into an explicit object with preconditions, steps, postconditions, recovery rules, and termination checks, so that it becomes executable, diagnosable, locally repairable, and consumable by another model.
We therefore propose ContractSkill, which converts a draft skill into an executable contract and improves it through deterministic verification, fault localization, and minimal patch repair. Starting from a model-generated draft, ContractSkill compiles the skill into an explicit artifact, executes it under a deterministic verifier, localizes failure to a specific step and error type, and edits only the implicated selector, condition, recovery rule, or action argument. Instead of treating failure as a reason to regenerate the whole skill, the framework treats it as a local repair problem over a persistent procedural object. The repaired artifact can then be reused by the source model or consumed by a target model without regenerating the skill.
We test this claim from three complementary angles. First, on VisualWebArena, ContractSkill improves execution in realistic multimodal web interaction where naive self-generated skills fail. Second, in MiniWoB, controlled comparisons show that the gain comes from deterministic verification, fault localization, and minimal repair rather than rewriting alone. Third, under matched transfer layers, repaired artifacts remain reusable across GLM-4.6V and Qwen3.5-Plus after removing the source model from the loop. This is a benchmark-specific test of portability rather than a claim of full-benchmark generalization.
@@
  • We identify the failure of self-generated web skills as a representation problem rather than only a generation problem, motivated by prior evidence that self-generated skills often fail to deliver the gains achieved by curated procedural knowledge.
  • We propose ContractSkill, an object-centered repair framework that converts a draft skill into an executable contract and improves the single skill artifact itself through deterministic verification, fault localization, and minimal patch operators.
  • We show that ContractSkill is effective on VisualWebArena, while MiniWoB validates the mechanism under controlled conditions. We further show that matched transfer layers provide evidence that repaired artifacts remain reusable across models after removing the source model from the loop. This is a benchmark-specific portability claim rather than one of full-benchmark generalization.
Figure 1 summarizes this artifact-centric workflow from draft generation to verifier-guided repair and downstream reuse.
@@ 2. Related Work
We organize prior work around the reliability of self-generated web skills, focusing on why stronger skill generation alone does not guarantee executable and reusable skill artifacts, how prior work improves behavior online or externalizes procedure to narrow this gap, and what verifier-guided evidence is needed to make such artifacts auditable and repairable under a fixed execution protocol.
2.1. Multimodal Web Agents
Recent benchmarks have rapidly expanded the scope of multimodal web agents. World of Bits and MiniWoB established controlled browser settings for basic interaction, while WebArena and VisualWebArena pushed evaluation toward longer-horizon, more realistic, and visually grounded web workflows (Shi et al., 2017; Liu et al., 2018; Zhou et al., 2024; Koh et al., 2024). Other resources, including Mind2Web, WebLINX, BrowserGym, WorkArena++, SeeAct, AutoWebGLM, AgentOccam, and scaling-oriented agent training efforts, further broaden the empirical landscape (Deng et al., 2023; Lu et al., 2024; de Chezelles et al., 2025; Boisvert et al., 2024; Zheng et al., 2024; Lai et al., 2024; Yang et al., 2025; Trabucco et al., 2025). These benchmarks are crucial for exposing realistic failure modes, but they mostly evaluate end-to-end agent behavior rather than how a self-generated web skill should be represented, verified, and repaired after execution breaks.
2.2. Self-Repair and Reflection in Agents
Within these environments, self-repair and reflection methods improve agent behavior by revising trajectories, reasoning traces, memories, or policies after observing failure (Shinn et al., 2023; Wang et al., 2024b; Yao et al., 2026; Weng et al., 2026; Yao et al., 2023; Zhao et al., 2024). These literatures are highly relevant for web agents because they shows that feedback-driven refinement can materially improve performance in interactive environments. However, the object being updated is usually the agent’s ongoing behavior or internal state in the current episode. By contrast, we repair an externalized skill artifact itself so that the result persists as a reusable object across episodes and can also be consumed by another model.
@@ 2.3. Skill Artifacts and Procedural Knowledge
A complementary line of work treats skills as explicit, reusable procedural knowledge rather than transient prompting context (Xu and Yan, 2026; Li et al., 2026; Li, 2026). Agent Workflow Memory stores reusable textual workflows, ASI induces and verifies executable programmatic skills online, PolySkill separates abstract goals from concrete implementations for cross-site reuse, SkillWeaver discovers website-specific APIs, Hierarchical Memory Tree organizes web experience with explicit pre/post conditions, and RAG-GUI retrieves external GUI guidance at inference time (Wang et al., 2025b, a; Yu et al., 2025; Zheng et al., 2025; Tan et al., 2026; Xu et al., 2025). This line is the closest to ours, but most prior work studies reusable knowledge at the level of workflow collections, APIs, or broader memory structures. Our focus is narrower and more operational: a single model-written multimodal web skill that must become executable, diagnosable, minimally repairable, and reusable across models.
@@ 2.4. Program Repair and Verifier-Guided Optimization
Once the skill itself becomes the repair target, program repair provides the methodological lens for our framework. Classic work such as GenProg, PAR, SemFix, DirectFix, Angelix, and Prophet emphasizes three ideas that are equally central here: localize the failure, constrain the edit space, and use a verifier to decide whether a patch should be accepted (Goues et al., 2012; Kim et al., 2013; Nguyen et al., 2013; Mechtaev et al., 2015, 2016; Long and Rinard, 2016). Verifier-guided refinement frameworks such as counterexample-guided abstraction refinement further show how failed checks can drive iterative improvement (Clarke et al., 2000). We borrow these principles not to repair source code directly, but to repair executable web-skill artifacts whose correctness is grounded in observable page-state evidence.
@@ 3.1. Environment
We study a multimodal web environment similar to VisualWebArena, WorkArena, and OSWorld (Koh et al., 2024; Drouin et al., 2024; Xie et al., 2024). An agent receives a natural-language task instruction $x$. At step $t$, the environment provides an observation $o_{t}=(I_{t},D_{t})$, where $I_{t}$ is the current page screenshot and $D_{t}$ is a structured page summary, such as a Document Object Model (DOM) abstraction, accessibility summary, or textual page state description. The agent outputs an action $a_{t}\in\mathcal{A}$, where $\mathcal{A}$ includes operations such as click, type, scroll, and select. The environment then transitions to a new observation $o_{t+1}$.
@@ 4.1. Skill Artifact Schema
ContractSkill represents each skill as an explicit artifact with the fields skill_name, goal, preconditions, steps, postconditions, recovery, and terminate. Each step contains four mandatory elements: selector, action, args, and post_assertion. Compared with free-form text, this schema exposes the exact operational assumptions behind the skill. The schema is designed to externalize procedures as reusable objects, which is also the motivating direction behind manual construction and page-level knowledge extraction in recent agent systems.
@@ 4.2. Deterministic Verifier
where $i$ is the failed step index, $e$ is a structured error code, and $\tau_{i}$ is the local trace evidence used for diagnosis. Verification uses deterministic checks such as URL matching, DOM retrieval, textual assertions, form-value checks, and state-change predicates. In practice, selectors are resolved on the current accessibility or DOM snapshot using role, visible text, and stable attributes rather than persistent node ids. Postconditions are instantiated from action-specific templates, such as value matching for type, state or navigation change for click, and option validation for select. This design keeps the repair signal grounded in observable page state rather than in model self-evaluation, matching the execution-oriented evaluation logic used in WebArena, VisualWebArena, and WorkArena (Zhou et al., 2024; Koh et al., 2024; Drouin et al., 2024). Table 1 summarizes the structured error codes returned by the verifier.
@@ 4.4. Minimal Patch Operators
Each operator changes only the artifact fragment tied to the diagnosed failure. This design preserves the rest of the artifact and avoids the instability that often accompanies full-text rewrites. The operator set intentionally favors small, interpretable edits, echoing the minimality bias emphasized in DirectFix and learning-guided repair systems such as Prophet (Mechtaev et al., 2015; Long and Rinard, 2016). Table 2 summarizes these minimal patch operators.
@@ 4.5. Verifier-Guided Repair Loop
(9) $\mathrm{Accept}(s^{\prime},s)=\begin{cases}1,&\text{if }J(s^{\prime})>J(s),\\ 0,&\text{otherwise.}\end{cases}$
@@ 5. Experimental Setup
Our experiments are organized around three questions. We ask (1) whether ContractSkill improves end-to-end web execution over direct acting and naive skill use, (2) whether any improvement comes specifically from verifier-guided local repair rather than rewriting alone, and (3) whether repaired skill artifacts remain reusable by another model under matched transfer layers after the source model is removed from the loop. Across all experiments, we treat efficiency as a success-cost trade-off rather than strict dominance over every baseline: the point is to test whether modest extra execution cost can buy substantially higher success when brittle self-generated skills fail. We structure the evaluation accordingly.
@@ 5.3. Baselines
We compare against two core baselines in all main-result tables. No-Skill executes the task directly without any external procedural guidance, representing the standard direct-acting setting. Self-Generated Skill lets the acting model produce a skill and immediately use it, but does not perform verifier-guided repair after failure. It therefore isolates the effect of skill writing alone, without the additional benefit of explicit verification or repair. On MiniWoB, we additionally include more controlled comparisons. Text-Only Rewrite rewrites the skill after failure in free-form text, without contracts, structured error codes, or localized edits, and therefore isolates the effect of rewriting alone. We also report two ablations of ContractSkill. No Failure Localization removes targeted identification of the failing skill component, and Unconstrained Repair permits broad revisions rather than restricted local fixes. These controlled variants help determine whether the gains of ContractSkill arise from structured verifier-guided repair.
@@ 5.4. Metrics
Our primary metric is verifier-based success rate. Under this framing, we also report average steps and average total tokens whenever those fields are present in the cleaned summaries. A repaired artifact may spend extra steps or tokens to complete tasks that a brittle self-generated skill fails to finish. Tokens are reported as average total tokens per task and abbreviated in thousands (k). For transfer, we compare repaired-artifact execution against the target model’s own self-generated skill baseline and report both absolute success and the point improvement.
@@ 6. Main Results
We report the main results in the order of argumentative strength. We first test whether ContractSkill works in realistic web interaction, then examine the same claim in a controlled environment where the mechanism is easier to isolate, and finally test whether the repaired artifact remains reusable under matched transfer layers after the source model is removed from the loop.
@@ 6.1. VisualWebArena
Table 3 reports the main VWA results. The overall pattern is clear. On realistic web tasks, simply having a skill is not enough, but repairing it into a contracted artifact is consistently beneficial. With GLM-4.6V, ContractSkill reaches 28.1% success, compared with 12.5% for No-Skill and 9.4% for Self-Generated Skill. The efficiency profile also improves. Average steps drop from 12.59 for No-Skill to 3.62, and average token usage decreases from 52.7k to 46.45k. With Qwen3.5-Plus, the packaged final verified aggregation raises ContractSkill to 37.5%, exceeding both No-Skill (28.1%) and Self-Generated Skill (10.9%), while keeping average token usage slightly below No-Skill (40.1k vs. 42.6k). On VWA, self-generated skills can underperform direct acting because an unverified but committed procedure often fails early and suppresses corrective exploration. As a result, brittle skills may use fewer steps and tokens precisely because they commit to the wrong procedure too soon. These results matter because VWA is precisely the regime in which brittle skills tend to fail, with long trajectories, changing page state, and multimodal ambiguity. The fact that ContractSkill improves both models suggests that the gain comes from structured repair of the external artifact rather than from model-specific prompting tricks.